OpenStudy (anonymous):

What degree measure is equal to 7π radians?

OpenStudy (lgbasallote):

you can do this the easy way...or the hard way... i suggest remembering that \(\pi = 180^o\)

OpenStudy (anonymous):

so it's 7*180?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

thanx, man

OpenStudy (lgbasallote):

welcome...btw..for future reference...to change radian into degree just multiply by \(\frac{180}{\pi}\) for example: \[ 7 \pi \times \frac{180}{\pi} \implies 7 \times 180\] same siz

OpenStudy (anonymous):

which is 1,260 degrees. right?

OpenStudy (lgbasallote):

that's right